# The Authentication Gap: Higher Education’s Widespread Noncompliance with NIST Digital Identity Guidelines

### v1

@@ Abstract
We examine the authentication practices of a diverse set of 101 colleges and universities in the United States and Canada to determine compliance with five standards in NIST Special Publication 800-63-3 Digital Identity Guidelines. We find widespread noncompliance with standards for password expiration, password composition rules, and knowledge-based authentication. Many institutions still require or recommend noncompliant practices despite years of expert advice and standards to the contrary. Furthermore, we observe that regional and liberal arts colleges have generally lower documented compliance rates than national and global universities, motivating further investment in authentication security at these institutions. These results are a wake-up call that expert cybersecurity recommendations are not sufficiently influencing the policies of higher education institutions, leaving the sector vulnerable to increasingly prevalent ransomware and other cyberattacks.
@@ 1 Introduction
Modern institutions are under constant threat of cyberattacks, including ransomware and phishing, that can completely cripple day-to-day operations, compromise sensitive data, and bear a hefty recovery cost [35]. Higher education institutions are especially vulnerable to these attacks as they process sensitive information and often lack resources to secure their infrastructure. In recent years, colleges and universities have been targeted by an increasing number of attacks [12]. A large portion of these attacks attempt to compromise authentication information of faculty, staff, or students in order to access and control sensitive data or resources.
To help mitigate potential attacks, the US National Institute of Standards and Technology (NIST) has published technical standards for digital authentication based on expert knowledge from the cybersecurity community [22]. Despite the availability and specificity of these standards and their updates over successive years, many institutions have been slow to adopt them in practice [18]. In 2022 and 2023, Lee et al. [27] and Hall et al. [24] found particularly low compliance with updated NIST authentication standards across the tech industry.
@@
In this paper, we investigate compliance of higher education institutions with NIST Special Publication (SP) 800-63-3 Digital Identity Guidelines. We examine the authentication practices of a diverse set of 101 colleges and universities in the United States and Canada, including public, private, R1, liberal arts, and regional institutions. We review online documentation of these institutions’ authentication practices and check for descriptions of and compliance with five key standards in NIST SP 800-63-3: 1) password expiration, 2) password composition guidance, 3) password strength meters, 4) knowledge-based authentication (security questions), and 5) multi-factor authentication. The recommended best practices for each of these standards have changed since previous versions of the standard, making them ideal candidates for identifying noncompliance.
Among institutions with authentication policy details available online, we find an unfortunately large fraction are out of compliance with the NIST standard. Contrary to NIST’s guidelines, password expiration, knowledge-based authentication, and specific password composition rules are still widespread. The rates of noncompliance we observe are higher than in Hall et al. [24]’s 2023 study of websites across the tech industry, indicating that higher education is out of step with the broader online ecosystem on this issue. We do, however, observe widespread deployment of NIST-recommended multi-factor authentication (MFA) across the institutions we examine. This finding is in keeping with recent literature about engagement with and opinions of MFA at educational institutions [10, 4, 16].
Importantly, we note that noncompliance rates are not distributed evenly across institution categories. US liberal arts colleges, US national universities, and Canadian global universities have more online documentation of digital authentication policies than US regional colleges. Across institutions with online documentation, US regional colleges generally have lower compliance rates than US liberal arts colleges, which have generally lower compliance rates than US national and Canadian global universities. Although our data do not provide a causal explanation for this result, it is reasonable to assume that lower compliance rates of US regional colleges are related to their generally more limited resources on an overall and per-student basis.
Our results do provide direct evidence that investments to strengthen the US and Canadian infrastructure against cyberattacks would be well directed towards improving authentication security in the higher education sector, especially at less well-resourced institutions. While out of scope for this study, we hope that follow-up research will examine digital authentication security at additional institutional categories, including community colleges, historically Black colleges and universities (HBCUs), and Hispanic-serving institutions (HSIs).
@@ 2 Related Work
The rapid rise in computational capabilities and credential leaks that allow attackers to break a large number of passwords have forced researchers and practitioners to re-examine authentication policies, including forced password expiration, password composition rules, knowledge-based authentication, strength meters, and multi-factor authentication. Scholars have also examined the actual behavior of individuals and companies regarding these authentication practices; however, no prior study has compared authentication policies of higher education institutions to current industry standards as broadly as in this work.
@@
Several prior efforts provided early evidence about the relative ineffectiveness of privacy expiration policies. In 2010, Zhang et al. [44] developed a framework that deduces new user passwords from old user passwords through a series of successive transformations. The framework was able to infer 41% of new passwords in an offline attack and 17% in an online attack. In 2014, Choong et al. [9] conducted a NIST study of US government employees’ password habits and found that when asked to create a new password, respondents tended to use less secure strategies, such as recycling old passwords or only making a minor change to an existing password.
In 2015, Farcasin and Chan-tin [17] surveyed university affiliates regarding pre-generated and expiring passwords. Respondents reported that a 120-day expiration time was too short, and the authors concluded that rapid expiration is untenable for most users, leading to password reuse and the creation of less secure new passwords. In 2018, Habib et al. [23] also surveyed users and found that regular password replacement usually led to similarly secure new passwords. They question the security gains of an expiration policy and recommend investing into alternative security measures.
In 2023, Gerlitz et al. [20] conducted a longitudinal study of employees of three German companies about the German Federal Office for Information Security’s removal of the password expiration requirement from their policy guidelines. While they reported a downward trend in requests to renew passwords, the investigation also revealed several factors that led to continued reliance of organizations on privacy expiration despite the federal recommendation against it. Several still, mistakenly, viewed the practice of periodically renewing passwords as beneficial to overall IT security. Several also kept password expiration while transitioning to MFA or were in industries with contradictory requirements, such as finance, which continue to mandate password expiration.
Knowledge-based Authentication.
Problems with knowledge-based authentication have been raised by the academic community for more than a decade. In 2009, Just and Aspinall [25] found that answers to security questions were typically low entropy (easier to crack) and users had trouble remembering their answers. In 2015, Bonneau et al. [6] examined personal knowledge questions at Google and found “a security level that is far lower than user-chosen passwords” with many users providing difficult-to-remember fake answers, concluding that “best practice should favor more reliable alternatives.”
@@
Multi-factor Authentication.
@@
Relatively few studies have examined compliance with the 2017 NIST SP 800-63-3 Digital Identity Guidelines. In 2023, Hall et al. [24] analyzed over 100 websites across industries that “report the most breaches in the Verizon Data Breach Investigation Report.” They found a mixture of compliance and noncompliance, including “nearly all websites…avoiding the use of security questions and SMS-based 2FA” (a substantially higher compliance rate than the institutions in our study), but that “many websites (greater than 80 percent) still deem ‘P@ssw0rd’ an acceptable password.” In 2022, Lee et al. [27] “examined the [password] policies of 120 of the most popular websites” and “found that only 13% of websites followed all relevant best practices…75% of websites do not stop users from choosing the most common passwords…45% burden users by requiring specific character classes in their passwords for minimal security benefit” and “low [19%] adoption of password strength meters.” Our work is the first to examine the compliance of higher education institutions with NIST Digital Identity Guidelines specifically and at scale.
3 NIST SP 800-63-3 Background and
Research Questions
NIST Special Publication 800-63-3 Digital Identity Guidelines [22] was published in 2017 as an update to SP 800-63-2 (2013). The SP provides “an overview of identity frameworks; using authenticators, credentials, and assertions in a digital system; and a risk-based process to select assurance levels.” The SP was developed in collaboration with the community, during which it received over 1,400 comments in a draft period from 2016–2017. The digital authentication standards presented in NIST SP 800-63-3 are extensive and cover practices across the technical stack and of differing relevance to different types of organizations.
We selected five standards from SP 800-63-3 to investigate for this study. We chose these specific standards because they
  1. 1.
    Concern key aspects of the user-facing digital authentication process
  2. 2.
    Are widely accepted by the technical security community
  3. 3.
    Require relatively little expense to implement
  4. 4.
    Have changed or been introduced since previous versions of the SP
Factors 1–3 above mean that all institutions should be aware of and have implemented these standards since the publication of SP 800-63-3. Factor 4 means that these standards are useful checks for institutional responsiveness to technical guidelines, because practices following pre-2017 versions of the standard will be readily detectable.
3.1 Password Expiration
In contrast to previous standards that online services should have their users change their passwords periodically, NIST SP 800-63B Section 5.1.1.2 paragraph 9 clearly states
Users tend to choose weaker memorized secrets when they know that they will have to change them in the near future. When those changes do occur, they often select a secret that is similar to their old memorized secret by applying a set of common transformations such as increasing a number in the password. This practice provides a false sense of security if any of the previous secrets has been compromised since attackers can apply these same common transformations. But if there is evidence that the memorized secret has been compromised, such as by a breach of the verifier’s hashed password database or observed fraudulent activity, subscribers should be required to change their memorized secrets. However, this event-based change should occur rarely, so that they are less motivated to choose a weak secret with the knowledge that it will only be used for a limited period of time. [32]
We investigate whether higher education institutions require or recommend regular password expiration or cycling for their affiliates.
3.2 Knowledge-based Authentication
The NIST SP 800-63 FAQ states
As noted in this quote, the use of KBA for digital authentication was formerly permitted. The practice is still widespread. Most laypeople are familiar with the need to provide answers to questions such as “What was the model of your first car?” “What was your grandmother’s maiden name?” and “What is your favorite winter sport?” (as well as a wide variety of others) during account creation and authentication. The use of such questions is no longer acceptable as noted above and any such use would be considered out of compliance.
We investigate whether higher education institutions require or recommend KBA during account creation, login, and/or recovery.
3.3 Password Composition Rules
NIST SP 800-63B Section 5.1.1.2 states that
This is in contrast to previous standards that online services require specific composition rules during password creation. It is approaching common knowledge that all passwords “should” contain at least one symbol, number, and (often) mix of capital and lowercase letters. Most are familiar with password creation interfaces that enforce specific composition rules. However, these interfaces are all out of compliance with the current NIST standard.
NIST SP 800-63B Appendix A.3 provides the rationale for this change:
We investigate whether higher education institutions require or recommend specific composition rules of this sort as a part of their password creation process.
3.4 Password Strength Meters
NIST SP 800-63B Section 5.1.1.2 states
Verifiers SHOULD offer guidance to the subscriber, such as a password-strength meter, to assist the user in choosing a strong memorized secret. This is particularly important…as it discourages trivial modification of listed (and likely very weak) memorized secrets [21]
NIST recognizes the need for authentication systems to help users create strong passwords and refers to strength meters as an example. Password strength meters are relatively common across online services, but have not received the same broad awareness as strict password composition rules. This is unfortunate, as password strength meters can be configured to account for a much wider range of weaknesses [13] that strict composition rules cannot prevent. Password strength meters are also more easily interpretable for users, often deploying colors and other interface elements to simplify guidance about a range of password vulnerabilities. However, password meters can be poorly configured, designed in a fashion that misinforms users, and may need to be supplemented with other forms of guidance. For example, some existing password meters may give the “green light” to weak passwords, such as “Jeremy1!” [18, 13].
We investigate whether higher education institutions provide password strength meters for their affiliates and require or recommend minimum strength values during password creation.
3.5 Multi-factor Authentication
According to NIST SP 800-63B Section 4.2, in order to achieve “authenticator assurance level 2” (“high confidence”) or stronger,
Multi-factor authentication (MFA) is well understood to be the best practice for digital authentication. Requiring two different forms of authentication, especially the combination of a memorized secret (e.g., a password) and a possession-based authenticator (i.e., “something you have”), makes account compromise significantly more challenging for would-be adversaries.
However, some authentication methods do not provide sufficient security to be deemed appropriate for use as a factor for MFA. NIST SP 800-63B Section 5.1.3.1 states that
[Authentication] methods that do not prove possession of a specific device, such as voice-over-IP (VOIP) or email, SHALL NOT be used for out-of-band authentication. [21]
We therefore investigate whether or not the institutions examined in this paper 1) recommend and/or require MFA and 2) allow email as an authentication factor.
@@ 4 Methods
This section presents our methods, including which higher education institutions we chose to investigate and why, how we systematically reviewed the publicly available online policies of these institutions for details about the digital authentication practices of interest (Section 3), our protocol for contacting institutional representatives to confirm or update the collected data, and some limitations that may arise from the nuances of our method.
@@ 4.1 Institution Selection
Specifically, we selected 101 institutions from the U.S. News and World Report lists of “top” institutions from several categories as of September 2023. We chose the top 20 US national universities [39], top 20 US liberal arts colleges [40], top 15 Canadian global universities [41], and a selection of US regional colleges consisting of the top 10 in each of the North, South, Midwest, and West regions [42]. Since multiple institutions may tie for a ranking on this list, this selection process resulted in 101 total institutions. The complete list of institutions is provided in Table 1.
US National Universities US Liberal Arts Colleges US North and South Regional Colleges US Midwest and West Regional Colleges Canadian Global Universities
Princeton University Williams College United States Coast Guard Academy Illinois Wesleyan University University of Toronto
Massachusetts Institute of Technology Amherst College Cooper Union for the Advancement of Science and Art Ohio Northern University University of British Columbia
Harvard University United States Naval Academy United States Merchant Marine Academy College of the Ozarks McGill University
Stanford University Pomona College Grove City College Taylor University University of Alberta
Yale University Swarthmore College Maine Maritime Academy Simpson College McMaster University
University of Pennsylvania Wellesley College Pennsylvania College of Technology Cottey College Universite de Montreal
@@
California Institute of Technology United States Air Force Academy Elmira College Alma College University of Calgary
@@
Duke University United States Military Academy at West Point Alfred State College–SUNY Benedictine College University of Waterloo
Brown University Bowdoin College College of Mount St. Vincent William Jewell College University of Ottawa
Johns Hopkins University Carleton College SUNY College of Technology at Canton Hiram College Western University
Northwestern University Barnard College University of Maine at Farmington Lake Superior State University Dalhousie University
Columbia University Claremont McKenna College High Point University Embry-Riddle Aeronautical University–Prescott Simon Fraser University
Cornell University Grinnell College Florida Polytechnic University California State University–Maritime Academy University of Victoria
University of Chicago Middlebury College Beacon College Carroll College University of Manitoba
University of California, Berkeley Wesleyan University Flagler College Criswell College Laval University
@@
University of California, Los Angeles Davidson College Wesleyan College Oregon Institute of Technology
Rice University Hamilton College Catawba College Fashion Institute of Design & Merchandising
Dartmouth College Harvey Mudd College University of the Ozarks College of Idaho
@@
Vanderbilt University Smith College Spring Hill College Brigham Young University–Hawaii
@@
University of Notre Dame Vassar College Huntingdon College Brigham Young University–Idaho
Colgate University Barton College University of Antelope Valley
Haverford College Newberry College
Washington and Lee University
Table 1: Higher education institutions included in this study. These institutions were selected from the 2023 U.S. News and World Report lists of top institutions in their respective categories [39, 40, 42, 41].
We recognize that this particular list of institutions is not without limitations, which we discuss in greater detail in Section 4.5. We strongly encourage follow-up studies investigating the digital authentication policies of different institutions, e.g., public flagship universities, historically Black colleges and universities (HBCUs), Hispanic-serving institutions (HSIs), and community colleges.

### v3

# Measuring NIST Authentication Standards Compliance by Higher Education Institutions

@@ Abstract
Technical standards are a longstanding method of communicating best practice recommendations based on expert consensus. Cybersecurity standards are particularly important for informing policies that protect critical systems and sensitive data. Measuring standards compliance is therefore essential to identify vulnerabilities arising from outdated policies and to determine whether expert advice has effectively diffused to practitioners. In this paper, we examine the authentication policies of a diverse set of 135 colleges and universities in the United States and Canada to determine compliance with four standards from NIST Special Publication 800-63 Digital Identity Guidelines. We find widespread, but not universal, deployment of multi-factor authentication across institutions. We also find prevalent outdated use of password expiration, password composition rules, and knowledge-based authentication. These results support further investment and research into incentive structures for standards compliance and the diffusion of expert guidance to practitioners.
[…]
@@ 1 Introduction
The 2004 National Institute of Standards and Technology (NIST) Special Publication (SP) 800-63 Electronic Authentication Guideline [15] advised users to secure their accounts with complex passwords composed of random characters, capital letters, and numbers, and to change their passwords regularly [53]. Federal agencies, corporations, and universities largely followed this advice by enforcing password complexity rules and regular password expiration. However, the expert consensus on authentication has changed dramatically in the intervening years as the field has broadened beyond computational cybersecurity to include human factors, usability, and other knowledge sources [31]. Recognizing flaws and limitations of the 2004 publication, NIST published major updates in 2011 [14], 2013 [13], and 2017 [36, 34, 35, 37]. Some of these updates reflected advances in authentication technologies; for example, multi-factor authentication is now strongly recommended. Others represented 180-degree changes from previous standards based on evolving expert knowledge; for example, forced password composition rules, regular password expiration, and knowledge-based authentication are now strongly discouraged [35].
For several reasons, these specific updates to NIST’s authentication standards, codified in the 2017 NIST SP 800-63 Digital Identity Guidelines [36, 34, 35, 37], create an ideal case study for examining how well expert cybersecurity knowledge is reaching practitioners responsible for system implementation, maintenance, and governance. First, the stark differences between the current and previous standards make it easy to detect stale institutional policies. Second, the current standards have been in place since 2017, meaning that all institutions have had ample time to update their policies if they are going to do so. Finally, the practices recommended by the current standards are widely endorsed by academic research (Section 3) and by other industry and government bodies [29, 43, 62, 18]. This means that noncompliance with these standards is more likely indicative of a general disconnect between expert recommendations and institutional policies rather than an inapplicability of the standards to a particular institution.
@@
In this paper, we examine the authentication policies of a diverse set of 135 colleges and universities in the United States and Canada, including public, private, R1, liberal arts, historically Black, and regional institutions. We review these institutions’ online documentation and check for compliance with the four standards from NIST SP 800-63 mentioned above regarding (1) multi-factor authentication, (2) password expiration, (3) password composition rules, and (4) knowledge-based authentication. We observe widespread, but not universal, deployment of multi-factor authentication across the institutions we examine. We also observe that password expiration, specific password composition rules, and knowledge-based authentication are still common contrary to NIST’s guidelines.
Prior work by Lee et al. [48] and Hall et al. [40] has found low compliance with updated NIST authentication standards across the tech industry; however, ours is the first study to date that has examined compliance by higher education institutions specifically. Focused attention on this sector is relevant, because higher education institutions have been increasingly targeted by cyberattacks attempting to compromise authentication credentials [28]. The widespread deployment of multi-factor authentication we observe is highly encouraging given that multi-factor authentication substantially increases the difficulty of credential-based attacks. Nevertheless, the outdated password and knowledge-based authentication policies we observe suggest a need for further examination of incentive structures and more research into knowledge diffusion barriers that result in noncompliant and insecure policies.
The rest of this paper is structured as follows. Section 2 details the specific standards from NIST SP 800-63 that we focus on in this paper, including NIST’s rationale for updating them from pre-2017 versions. Section 3 reviews related academic research supporting the standards updates, measuring compliance with NIST SP 800-63 in other contexts, and about authentication practices in higher education generally. Section 4 describes our data collection and analysis methods, including research ethics and limitations. Section 5 presents our results, including rates of compliance and noncompliance across institutional categories. Section 6 discusses the implications of these results, emphasizing their significance for the usable security community. Section 7 concludes.
2 Background and Research Questions
NIST SP 800-63 Digital Identity Guidelines was published in 2017 as a four-volume document suite (SP 800-63-3 [36], SP 800-63A [34], SP 800-63B [35], and SP 800-63C [37]). This was an update to SP 800-63-2 (2013) [13] and as part of a series of authentication standards documents including SP 800-63-1 (2011) [14] and SP 800-63 (2004) [15]. The 2017 version was developed in collaboration with the community, during which it received over 1,400 comments in a draft period from 2016–2017. NIST SP 800-63 provides “an overview of identity frameworks; using authenticators, credentials, and assertions in a digital system; and a risk-based process to select assurance levels” [36]. The digital authentication standards presented in NIST SP 800-63 are extensive and cover practices across the technical stack and of differing relevance to different types of organizations.
We selected four standards from NIST SP 800-63 to investigate for this study. We chose these specific standards because they form an ideal case study of the diffusion of expert cybersecurity knowledge to practitioners for reasons presented in Section 1 and summarized here: First, clear differences between the current and previous versions of these standards make it easy for us to detect outdated institutional policies. Second, these standards have been in place since 2017, providing institutions ample time to update their policies. Third, these standards are widely supported by academic research and other government bodies, meaning that noncompliance is more likely due to a general disconnect between expert recommendations and institutional policies than due to conflicting expert advice.
2.1 Multi-factor Authentication
Multi-factor authentication (MFA) is well understood to be the best practice for digital authentication. Requiring two or more different forms of authentication, especially the combination of a memorized secret (e.g., a password) and a possession-based authenticator (i.e., “something you have”), makes account compromise significantly more challenging for would-be adversaries.
Pre-2017 versions of NIST SP 800-63 used a "levels of assurance" (LOA) model to determine what authentication procedures should be followed. The required level depended on the amount of risk posed by an authentication error. Level 1 (no risk) did not require or recommend MFA. Level 2 (moderate risk) recommended, but did not require, MFA. Only Levels 3 and 4 (high and very high risk) required MFA.
The 2017 version of NIST SP 800-63 changed to a more nuanced model with separate levels for identity assurance (IAL), authenticator assurance (AAL), and federation assurance (FAL). AAL 1 (low risk) recommends, but does not require, MFA. AAL 2 (moderate risk) requires MFA:
AAL 3 (high risk) requires MFA with a hardware-based authenticator. The inclusion of a MFA requirement for situations where an authentication error poses “moderate risk” is a significant expansion of of the MFA requirement from the pre-2017 versions. Since authentication errors at higher education institutions pose at least “moderate risk,” we investigate whether or not the institutions examined in this paper require MFA.
2.2 Password Expiration
All pre-2017 versions of NIST SP 800-63 recommended regular password expiration or cycling as a best practice. This received a 180-degree update in the 2017 version, as stated in NIST SP 800-63B Section 5.1.1.2:
Users tend to choose weaker memorized secrets when they know that they will have to change them in the near future. When those changes do occur, they often select a secret that is similar to their old memorized secret by applying a set of common transformations such as increasing a number in the password. This practice provides a false sense of security if any of the previous secrets has been compromised since attackers can apply these same common transformations. [62]
We therefore investigate whether higher education institutions require or recommend regular password expiration or cycling for their affiliates.
2.3 Password Composition Rules
All pre-2017 versions of NIST SP 800-63 recommended password composition rules during account creation. The 2017 version of the standard makes a 180-degree change to this recommendation, as stated in NIST SP 800-63B Section 5.1.1.2:
It is approaching common knowledge that all passwords “should” contain at least one symbol, number, and (often) mix of capital and lowercase letters. Most are familiar with password creation interfaces that enforce specific composition rules. However, these interfaces are all out of compliance with the current NIST standard.
NIST SP 800-63B Appendix A.3 provides the rationale for this change:
We therefore investigate whether higher education institutions require or recommend specific composition rules as a part of their password creation processes.
2.4 Knowledge-based Authentication
All pre-2017 versions of NIST SP 800-63 recommended knowledge-based authentication (KBA) as a component of a strong authentication posture. The 2017 version makes a 180-degree change to this recommendation, as stated in the NIST SP 800-63 FAQ:
Despite this change, the use of KBA for digital authentication is still widespread. Most laypeople are familiar with the need to provide answers to questions such as “What was the model of your first car?” “What was your grandmother’s maiden name?” and “What is your favorite winter sport?” (as well as a wide variety of others) during account creation and authentication. The use of such questions is no longer acceptable. We therefore investigate whether higher education institutions require or recommend KBA during account creation, login, and/or recovery.
2.5 Canadian Standards
The NIST standards we investigate have direct analogues in Canadian standards. For example, the Password Guidance published by the Treasury Board of Canada Secretariat recommends that system owners “use two-factor authentication (2FA),” “eliminate password expiry,” and “disable or reduce complexity policies” [73]. The congruence of these standards supports our inclusion of 20 Canadian universities in this study.
@@ 3 Related Work
The rapid rise in computational capabilities and credential leaks that allow attackers to crack a large number of passwords have forced researchers and practitioners to re-examine authentication policies, including multi-factor authentication, forced password expiration, password composition rules, and knowledge-based authentication. Scholars have also examined the actual behavior of individuals and companies regarding these and other authentication practices [23, 47, 9, 54, 66, 64]; however, no prior study has compared authentication policies of higher education institutions to current industry standards as broadly as in this work.
Multi-factor Authentication.
@@
Several prior efforts provided early evidence about the relative ineffectiveness of password expiration policies. In 2010, Zhang et al. [82] developed a framework that deduces new user passwords from old user passwords through a series of successive transformations. The framework was able to infer 41% of new passwords in an offline attack and 17% in an online attack. In 2014, Choong et al. [16] conducted a NIST study of US government employees’ password habits and found that when asked to create a new password, respondents tended to use less secure strategies, such as recycling old passwords or only making a minor change to an existing password.
In 2015, Farcasin et al. [27] surveyed university affiliates regarding pre-generated and expiring passwords. Respondents reported that a 120-day expiration time was too short, and the authors concluded that rapid expiration is untenable for most users, leading to password reuse and the creation of less secure new passwords. In 2018, Habib et al. [39] also surveyed users and found that regular password replacement usually led to similarly secure new passwords. They question the security gains of an expiration policy and recommend investing into alternative security measures.
In 2023, Gerlitz et al. [33] conducted a longitudinal study of employees of three German companies about the German Federal Office for Information Security’s removal of the password expiration requirement from their policy guidelines. While they reported a downward trend in requests to renew passwords, the investigation also revealed several factors that led to continued reliance of organizations on privacy expiration despite the federal recommendation against it. Several organizations still, mistakenly, viewed the practice of periodically renewing passwords as beneficial to overall IT security. Several organizations also kept password expiration while transitioning to MFA or were in industries with contradictory requirements, such as finance, which continue to mandate password expiration.
@@
Authentication experts and NIST SP 800-63 now support the use of password strength meters and “increased password length…especially through encouraging the use of passphrases” [62] instead of forced composition rules. Meters and other forms of visual feedback are less likely to lead to formulaic passwords while still giving users an indication of password strength. However, the design of the feedback matters. In 2008, Forget et al. [30] proposed a system that added random characters inside user-chosen passwords to increase their entropy. In 2017, Segreti et al. [69] tested adaptive password composition policies that change over time as users create new passwords. In 2023, Behfar et al. [8] explored the effectiveness of various metaphor-based password strength indicators on strong user password selection. In 2023, Amador et al. [5] found that interventions guided by prospect theory can cause users to improve password strength. In 2024, Paudel et al. [63] designed and evaluated a “priming-through-persuasion” approach to inform users about weak password selections.
Knowledge-based Authentication.
Problems with knowledge-based authentication have been raised by the academic community for more than a decade. In 2009, Just et al. [42] found that answers to security questions were typically low entropy (easy to crack) and users had trouble remembering their answers. In 2015, Bonneau et al. [10] examined personal knowledge questions at Google and found “a security level that is far lower than user-chosen passwords” with many users providing difficult-to-remember fake answers, concluding that “best practice should favor more reliable alternatives.”
@@
Relatively few studies have examined compliance with the 2017 NIST SP 800-63 Digital Identity Guidelines. In 2022, Lee et al. [48] “examined the [password] policies of 120 of the most popular websites” and “found that only 13% of websites followed all relevant best practices…75% of websites do not stop users from choosing the most common passwords…45% burden users by requiring specific character classes in their passwords for minimal security benefit.” In 2023, Hall et al. [40] analyzed over 100 websites across industries that “report the most breaches in the Verizon Data Breach Investigation Report.” They found a mixture of compliance and noncompliance, including “nearly all websites…avoiding the use of security questions and SMS-based 2FA” (a higher compliance rate than the institutions in our study), but that “many websites (greater than 80 percent) still deem ‘P@ssw0rd’ an acceptable password.” Our work is the first to examine the compliance of higher education institutions with NIST SP 800-63 Digital Identity Guidelines specifically and at scale.
@@ 4 Methods
This section presents our methods, including which higher education institutions we chose to investigate and why, how we reviewed publicly available online documentation from these institutions for details about the authentication policies of interest (Section 2), and some limitations that may arise from the nuances of our methods.
@@ 4.1 Institution Selection
Specifically, we selected institutions from the U.S. News and World Report lists of “top” institutions from several categories as of September 2023. We chose the top 20 Canadian global universities [75], top 20 US national universities [76], top 20 US public colleges and universities [80], top 20 US liberal arts colleges [79], top 10 US historically Black colleges and universities (HBCUs) [78], and a selection of US regional colleges consisting of the top 10 in each of the North, South, Midwest, and West regions [^1] Less one institution that closed during the study (University of Antelope Valley). [77]. Since multiple institutions may tie for a ranking on these lists, this selection process resulted in 135 total institutions (Table 1). We recognize that this particular list of institutions is not without limitations, which we discuss in greater detail in Section 4.6.
US National Universities MFA PE CR KBA US Public Colleges and Universities MFA PE CR KBA US Liberal Arts Colleges MFA PE CR KBA
Princeton University A N X N University of Michigan–Ann Arbor A N N X Williams College A S A N
Massachusetts Institute of Technology A A A N University of North Carolina at Chapel Hill A A S A Amherst College A S A N
Harvard University A N R N University of Virginia A A A R United States Naval Academy N A A A
Stanford University A N S A University of California, Davis A N N A Pomona College A X A N
Yale University A A S X University of California, San Diego A N S A Swarthmore College A A R N
University of Pennsylvania A X S N University of Florida A A A N Wellesley College A R A N
@@
California Institute of Technology A N R R University of Texas at Austin A N A N United States Air Force Academy A N A N
@@
Duke University A X X N Georgia Institute of Technology A A A A United States Military Academy at West Point A N S N
Brown University A R A R University of California, Irvine A R A A Bowdoin College A N A A
Johns Hopkins University A A A A University of California, Santa Barbara A R R A Carleton College A S S N
Northwestern University A A A A University of Illinois Urbana-Champaign A A A N Barnard College A N N N
Columbia University A D A A University of Wisconsin–Madison A R R A Claremont McKenna College A R R S
Cornell University A N A D Rutgers University–New Brunswick A N N A Grinnell College S N S N
University of Chicago A S S A University of Washington A N A D Middlebury College A N A N
University of California, Berkeley A N A S The Ohio State University A A A A Wesleyan University A S N N
@@
University of California, Los Angeles A N R N Purdue University–Main Campus A A A A Davidson College R N X N
Rice University A R A N University of Maryland, College Park A A A X Hamilton College A X X A
Dartmouth College A D X X Texas A&M University S S A R Harvey Mudd College A N A R
Vanderbilt University A A N R University of Georgia A S A A Smith College A R N S
University of Notre Dame A X N A Virginia Tech A A S N Vassar College A N N N
Colgate University A N R N
Haverford College A N N N
Washington and Lee University A X A N
Canadian Global Universities MFA PE CR KBA US Historically Black Colleges and Universities MFA PE CR KBA US Regional Colleges – North MFA PE CR KBA
University of Toronto A N R R Spelman College R N A N United States Coast Guard Academy N N S N
University of British Columbia A A R A Howard University A N N A Cooper Union for the Advancement of Science and Art A R R N
McGill University A N N R Florida A&M University A A A A United States Merchant Marine Academy N A N N
University of Alberta S N A N Tuskegee University N R A N Grove City College N N N N
McMaster University A R A N Morehouse College A S N N Maine Maritime Academy S A A N
Universite de Montreal N N A N Xavier University of Louisiana A N A N Pennsylvania College of Technology S N N N
University of Calgary A N R N Hampton University S S N S Elmira College R N N N
University of Waterloo A R X S North Carolina A&T State University S R R N Alfred State College–SUNY S A A R
University of Ottawa A A A N Delaware State University N A A A College of Mount St. Vincent A N A N
Western University A R A S Morgan State University A A N A SUNY College of Technology at Canton S R R N
Dalhousie University A N X N University of Maine at Farmington R A A A
Simon Fraser University A A S N
University of Victoria A N N N
University of Manitoba A A A R
Laval University A N A N
York University - Canada A A R R
Queens University - Canada A A R A
University of Saskatchewan A N A R
University of Guelph A R A R
Carleton University A N A R
US Regional Colleges – South MFA PE CR KBA US Regional Colleges – Midwest MFA PE CR KBA US Regional Colleges – West MFA PE CR KBA
High Point University A N N N Illinois Wesleyan University A N A N Embry-Riddle Aeronautical University–Prescott S N R A
Florida Polytechnic University A A N N Ohio Northern University N N N N California State University–Maritime Academy A A A N
Beacon College S S S N College of the Ozarks N N N N Carroll College N N N N
Flagler College A S A A Taylor University S N A N Criswell College S N N N
Wesleyan College S N A A Simpson College N N A N Oregon Institute of Technology N A A A
Catawba College A N A A Cottey College N N N N Fashion Institute of Design & Merchandising N N N N
University of the Ozarks N N N N Alma College S A S N College of Idaho A A A N
@@
Spring Hill College N A N A Benedictine College A A A N Brigham Young University–Hawaii A N R A
@@
Huntingdon College N N N N William Jewell College R S A N Brigham Young University–Idaho A R N R
Barton College N R N N Hiram College A S A A
Newberry College N S A N Lake Superior State University N N N N
Table 1: Publicly available policies regarding the four authentication standards from NIST SP 800-63 investigated in this study: MFA = multifactor authentication, PE = password expiration, CR = password composition rules, KBA = knowledge-based authentication. Code abbreviations refer to the codebook in Section 4.2: A = required for all affiliates, S = required for specific affiliates, R = recommended, D = discouraged, X = disallowed, N = no information found. These institutions were selected from the 2023 U.S. News and World Report lists of top institutions in their respective categories [76, 79, 77, 75, 80, 78]. Additional information, including specific affiliates, password expiration frequencies, policy links, and policy screenshots or PDF documents are available at https://github.com/NoahApthorpe/apthorpe-soups-2025-data.
4.2 Codebook Preparation
Since each of the NIST standards we investigated has a “should” (“shall”) or “should not” format, we used a deductive coding approach with a fixed set of codes determined prior to data collection. We chose a codebook applicable to all of the standards that would be easily understood by trained undergraduate researchers. Each of the six codes were potential answers to the following question: “Based on publicly available online documentation, what is the institution’s policy regarding the practice associated with this standard?”
  1. 1.
    “Required”: The practice is required for all affiliates of the institution.
  2. 2.
    “Required for Specific Affiliates”: At least some affiliates of the institution are required to follow the practice. In this case, we also recorded which specific affiliates are subject to this requirement.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
    “No Information Found”: The institution’s publicly available online documentation does not provide any or enough information to determine the institution’s policy regarding the practice.
We confirmed the effectiveness of this codebook through trial data collection with three institutions. The codes covered all cases appearing in the trial.